I'm trying to help my grandson solve this problem.. only I don't remember how to do it.
Thank you for the help
Kid has a bug collection of insects(6legs) and arachnids(8legs) he has a total of 8 bugs. With a combined total of 54 legs. How many of each type of bug does he have. Write a formula and solve.

Found 2 solutions by Alan3354, richwmiller:
Answer by Alan3354(69443)

Kid has a bug collection of insects(6legs) and arachnids(8legs) he has a total of 8 bugs. With a combined total of 54 legs. How many of each type of bug does he have. Write a formula and solve.
----------------
I + A = 8 (Total of 8 bugs)
6I + 8A = 54 (Total legs)
---------------------------
I = 8 - A
Sub for I into 2nd eqn
6(8-A) + 8A = 54
48-6A + 8A = 54
A = 3
I = 5

Answer by richwmiller(17219)

This the same as a coin problem
nickels are worth 5 cents
Dimes are 10 cents
What do we know?
Insects have 6 legs
Spiders (arachnids) have 8 legs
There are a total of 54 legs
and there are a total of 8 bugs
a+i=8
6i+8a=54
a=3
i=5